OpenStudy (anonymous):

Divide 120 by 37 and give me a mixed number! I know this is basic but please help me understand it. If you can go step by step!

OpenStudy (saifoo.khan):

can u join me on twiddla?

OpenStudy (anonymous):

How many times does 120 go into 37?

OpenStudy (anonymous):

If you know that, then all you have to do is find the remainder.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

But finding the remainder and changing it into a mixed number is the hard part for me!

OpenStudy (anonymous):

Right. So the answer will be 3 and some fraction

OpenStudy (anonymous):

The remainder will be 120 - 3*37

OpenStudy (anonymous):

What do you get for that?

OpenStudy (anonymous):

35.608 or 35.61

OpenStudy (anonymous):

No no. What is 120 - 3 * 37

OpenStudy (anonymous):

not divided, times.

OpenStudy (anonymous):

116.63

OpenStudy (anonymous):

How do you get a decimal answer? 3*37 = 111 120 - 111 = 9

OpenStudy (anonymous):

So your remainder is 9

OpenStudy (anonymous):

Therefore the mixed number will be 3 and 9/37

OpenStudy (anonymous):

Sorry my fault i was think sub not multiplication!

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

So in general if you have n/d and you want it as a mixed number: \[\lfloor{n/d} \rfloor \text{ and } \frac{n - \lfloor{n/d} \rfloor}{d}\]

OpenStudy (anonymous):

where \(\lfloor{n/d} \rfloor\) is the whole number of times d goes into n.